You need to replace 3/4 mile of fence. You do not need to add the first corner pole. If you put a pole in every 1/12 of a mile, how many poles do you need? Please show your work.
1
Expert's answer
2019-01-11T13:11:42-0500

The target number N of poles is defined by the ratio of the total length of the fence to be replaced to the pole gap:

"N=\frac{3}{4}:\frac{1}{12}=\frac{3\cdot12}{4\cdot1}=9."